OpenStudy (anonymous):

a circular container is 1/3 full of water. after 20 m^3 of water was added into, the container is now 7/9 full. what is the volume of the container when full?

OpenStudy (lgbasallote):

here's a hint \[\frac 13 x + 20 = \frac 79 x\] where x = volume of container when full solve for x does that help?

OpenStudy (anonymous):

@lgbasallote does the answer is 44.4444?

OpenStudy (anonymous):

x = 45

OpenStudy (lgbasallote):

i don't think so...check again

OpenStudy (lgbasallote):

remember to change 1/3 into common denominator

OpenStudy (lgbasallote):

if you change 1/3 into common denominator you get \[\implies \frac 39 x + 20 = \frac 79 x\] try to solve for x again then tell me what you get @Simonbane